solve the system of equations below and enter your answer as an ordered pair.
4y=3x+5
6y=3x=21

Answer by mananth(16949)

-3 x + 4 y = 5 .............1
Total value
-3 x + 6 y = 21 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -1
-3 x 4 y = 5
3 x -6 y = -21
Add the two equations
-2y=-16
y=8
plug y in (1)
-3x+4y=5
-3x+4*8=5
-3x+32=5
-3x=-32+5
-3x=-27
x=9
(8,9) is the solution